# Path Gradients after Flow Matching

### v1

@@ Appendix A Appendix
$\mathcal{N}\left(\begin{bmatrix}a_{i}\\ b_{i}\end{bmatrix},\begin{bmatrix}c_{i}+0.01&0\\ 0&d_{i}+0.01\end{bmatrix}\right)\,,$ (22)
@@
Figure 5 shows the weighted difference
@@
Figure 5: Loss in space after training as done in Figure 1
@@ A.3 Datasets
[…]
@@ A.5 Additional results for alanine dipeptide
Figure 6: Alanine dipeptide results for the TBG model and the classical force field with and without Path Gradient finetuning. Left and middle: Free energy projection for the $\varphi$ and $\psi$ dihedral angles, respectively. Right: Bond-length distribution for the Carbon - Nitrogen bond.
Here we present additional alanine dipeptide results for the TBG model in Figure 6. We observe that the $\varphi$ and $\psi$ dihedral-angle projections change little after path-gradient fine-tuning. Moreover, reweighting to the target distribution via Equation 2 succeeds in both cases, though it is more efficient post–fine-tuning, as reflected by a higher ESS (see Table 1). In contrast, all bond-length distributions align much more closely with the target after fine-tuning, exemplified here by the Carbon - Nitrogen bond in Figure 6.
@@ A.6 Additional results for alanine dipeptide - classical with more memory
For classical AD2, we further investigated the performance of fine-tuning with PGs when allowed a larger memory footprint. We fine-tuned with batch-size 1024 for two epochs, which resulted in $22GB$ VRAM usage and a runtime of 92 minutes. In Table 2 we can see that this yielded a further improvement over Table 1.
@@
[…]
Only FM $-128.01\pm 0.01$ $14.42\pm 2.44$ $12.21\pm 1.06$
@@ A.7 Additional results for experiments on flow trajectory length
Fine-tuning on the 13-particle Lennard–Jones system used an A100 for 10 epochs (152 min/run) starting from the Klein et al. (2023b) checkpoints, with training and evaluation on $5\times 10^{5}$ samples. As shown in Table 3, path-gradient fine-tuning boosts NLL and ESS across the board while leaving flow trajectory lengths nearly unchanged (see Figure 7). One of the nine runs diverged during training, was discarded and repeated.
@@
Method Fine-tuning Trajectory Length NLL ESS-Q (%) ESS-P (%)
Standard before $3.76\pm 0.00$ $-16.02\pm 0.01$ $60.45\pm 0.89$ $40.23\pm 20.11$
after $3.86\pm 0.00$ $-16.22\pm 0.00$ $85.78\pm 0.69$ $85.93\pm 0.57$
Optimal Transport before $2.80\pm 0.01$ $-16.01\pm 0.01$ $57.16\pm 0.88$ $57.07\pm 0.54$
after $2.78\pm 0.04$ $-16.22\pm 0.00$ $85.91\pm 0.03$ $85.78\pm 0.02$
Equivariant Op- before $2.11\pm 0.00$ $-16.04\pm 0.00$ $58.11\pm 1.93$ $19.77\pm 12.36$
timal Transport after $2.19\pm 0.00$ $-16.23\pm 0.00$ $87.77\pm 0.24$ $58.40\pm 29.20$
@@
A.8 Code libraries
A.9 Pseudocode for forward KL path gradients via augmented adjoint
@@
2:     $\dot{x},\dot{\text{div}}\leftarrow\text{black\_box\_dynamics}(t,x_{t})$triangleright Implement Augmented Adjoint State ODE
@@
3:     $\nabla\log p_{1}\leftarrow\text{gradient}_{x_{1}}(\log p_{1})$
[…]
@@
6:     $\nabla\log q_{0}\leftarrow\text{gradient}_{x_{0}}(\log q_{0})$
7:     $\nabla_{x_{0}}\mathcal{L}\leftarrow\frac{1}{N}\left(\nabla\log p_{0,\theta}-\nabla\log q_{0}\right)$
8:     $\text{path gradients}\leftarrow\text{gradient}_{\theta}\left(x_{0}\cdot\texttt{detach}(\nabla_{x_{0}}\mathcal{L})]\right)$triangleright Backpropage using standard Adjoint state method
9:return path gradients
10:end function

### v2

@@ 5.2 Future Work
5.3 Broader Impact
This foundational research has no immediate societal impact, but if scalable, it could be used to accelerate drug and materials discovery by replacing MD simulations. Potential risks include misuse for biothreat development and the lack of convergence guarantees, which may lead to incomplete sampling and misleading conclusions.
Acknowledgements
LV wants to thank Pan Kessel and Shinichi Nakajima for the helpful discussions. LK gratefully acknowledges support by BIFOLD - Berlin Institute for the Foundations of Learning and Data.
@@ Appendix A Appendix
[…]
@@
Figure˜5 shows the weighted difference
@@
Figure 5: Loss in space after training as done in Figure˜1
@@ A.3 Datasets
The classical alanine dipeptide dataset was generated with an MD simulation, using the classical Amber ff99SBildn force-field at $300\textrm{K}$ for implicit solvent for a duration of $1$ ms Köhler et al. (2021) with the openMM library. The datasets is available as part of the public bgmol (MIT licence) repository here: https://github.com/noegroup/bgmol.
@@ A.5 Additional results for alanine dipeptide
Figure 6: Alanine dipeptide results for the TBG model and the classical force field with and without Path Gradient finetuning. Left and middle: Free energy projection for the $\varphi$ and $\psi$ dihedral angles, respectively. Right: Bond-length distribution for a Carbon - Nitrogen bond.
Here we present additional alanine dipeptide results for the TBG model in Figure˜6. We observe that the $\varphi$ and $\psi$ dihedral-angle projections change little after path-gradient fine-tuning. Moreover, reweighting to the target distribution via Equation˜2 succeeds in both cases, though it is more efficient post–fine-tuning, as reflected by a higher ESS (see Table˜1). In contrast, all bond-length distributions align much more closely with the target after fine-tuning, exemplified here by a Carbon - Nitrogen bond in Figure˜6.
@@ A.6 Additional results for alanine dipeptide - classical with more memory
For classical AD2, we further investigated the performance of fine-tuning with PGs when allowed a larger memory footprint. We fine-tuned with batch-size 1024 for two epochs, which resulted in $22GB$ VRAM usage and a runtime of 92 minutes. In Table˜2 we can see that this yielded a further improvement over Table˜1.
@@
Training type NLL $(\downarrow)$ ESSq $\%(\uparrow)$ ESSp $\%(\uparrow)$
Only FM $128.01\pm 0.01$ $14.42\pm 2.44$ $12.21\pm 1.06$
@@ A.7 Additional results for experiments on flow trajectory length
Fine-tuning on the 13-particle Lennard–Jones system used an A100 for 10 epochs (152 min/run) starting from the Klein et al. (2023b) checkpoints, with training and evaluation on $5\times 10^{5}$ samples. As shown in Table˜3, path-gradient fine-tuning boosts NLL and ESS across the board while leaving flow trajectory lengths nearly unchanged (see Figure˜7). One of the nine runs diverged during training, was discarded and repeated.
@@
Method Fine-tuning Trajectory Length NLL ESSq (%) ESSp (%)
Standard before $3.76\pm 0.00$ $16.02\pm 0.01$ $60.45\pm 0.89$ $40.23\pm 20.11$
after $3.86\pm 0.00$ $16.22\pm 0.00$ $85.78\pm 0.69$ $85.93\pm 0.57$
Optimal Transport before $2.80\pm 0.01$ $16.01\pm 0.01$ $57.16\pm 0.88$ $57.07\pm 0.54$
after $2.78\pm 0.04$ $16.22\pm 0.00$ $85.91\pm 0.03$ $85.78\pm 0.02$
Equivariant Op- before $2.11\pm 0.00$ $16.04\pm 0.00$ $58.11\pm 1.93$ $19.77\pm 12.36$
timal Transport after $2.19\pm 0.00$ $16.23\pm 0.00$ $87.77\pm 0.24$ $58.40\pm 29.20$
@@
A.8 Comparison with maximum likelihood training
We here present comparisons with maximum likelihood training as a baseline. To this end we compare to Garcia Satorras et al. (2021), which is nearly the same architecture as in Klein et al. (2023b), but the model is trained via maximum likelihood training instead of flow matching. The results are shown in Table˜4.
Method NLL $\textbf{ESS}_{q}$ in $\%$
ML training (Garcia Satorras et al., 2021) $15.83\pm 0.07$ $39.78\pm 6.19$
Only FM (Klein et al., 2023b) $16.09\pm 0.03$ $54.36\pm 5.43$
Hybrid approach (Ours) $16.21\pm 0.00$ $82.97\pm 0.40$
Table 4: Results for the Lennard-Jones system with $13$ particles (LJ13).
A.9 Additional energy histograms
We show additional energy histograms of the systems investigated in Section˜4 in Figure˜8. The average energy for every investigated system is smaller and closer to the target energy distribution after PG finetuning.
Figure 8: Energy histograms for the a) LJ13 systems, b) the LJ55 system, and c-i) Different didpeptides from the testset for the transferable Boltzmann Generator (TBG).
A.10 Results for transferable Boltzmann Generators (TBG) evaluated on dipeptides
We applied the hybrid approach to transferable Boltzmann Generators (TBG) on dipeptides as introduced in Klein & Noé (2024). We again fine-tune the pretrained Boltzmann Generator from Klein & Noé (2024) with path gradients. Training took 6 days on an A100 with learning rate 0.00001. This transferable Boltzmann Generator is trained on a subset of all possible dipeptides and evaluated on unseen ones (for more details see Section˜A.3). Due to the expensive evaluation, we evaluated the model on $16$ test dipeptides and chose the same subset as in Klein & Noé (2024). Our experiments show that fine-tuning with path gradients improves the NLL and energies for all evaluated unseen test dipeptides with an average improvement to -100.93 from -100.72 NLL and also shows average improvements for the ESSq of around 23% to an efficiency of $9.79\%$ as shown in Table˜5. An example energy histogram is shown in Figure˜8c.
Dipeptide NLL (Before) NLL (After PG) ESSq (Before) in $\%$ ESSq (After PG) in $\%$
KS -100.82 -100.99 6.0 4.6
AT -75.49 -75.61 20.8 21.5
GN -65.45 -65.55 19.2 25.6
LW -148.91 -149.22 0.4 3.6
NY -118.24 -118.47 9.5 10.5
IM -106.80 -107.01 3.1 4.6
TD -81.09 -81.19 4.1 11.6
HT -103.06 -103.27 0.2 5.5
KG -88.60 -88.79 5.0 7.4
NF -116.45 -116.67 3.7 12.2
RL -135.71 -136.12 1.3 1.5
ET -89.94 -90.08 6.3 4.1
AC -63.90 -63.91 31.8 33.3
GP -71.63 -71.77 10.5 8.0
KQ -119.36 -119.68 4.3 2.0
RV -126.28 -126.62 1.3 0.6
Average -100.73 -100.93 7.96 9.79
Table 5: Comparison of NLL and ESS before and after PG for different dipeptides from the testset.
A.11 About the different estimators
In the following, we first discuss the Maximum Likelihood (ML) and Path Gradient (PG) estimators for the forward KL and show benefits of PG over FM for a toy example.
A.11.1 The Maximum Likelihood and Path Gradient Estimators
In short: both the ML and PG estimators are unbiased and consistent, but they differ in variance. For PG, we can give guarantees about their variance, once $q_{\theta}$ equals to $p$, see e.g. (Vaitl, 2024; Roeder et al., 2017; Tucker et al., 2019; Vaitl et al., 2022b).
Both Maximum Likelihood and the Path Gradient estimators optimize the Forward KL Equation 8. Let’s look at both estimators in detail and compare them. A full derivation can be found in Appendix B.3.2 of Vaitl et al. (2024) & Chapter 4 of Vaitl (2024).
To obtain the estimator $\mathcal{G}_{ML}$, we first observe that the first term in the KL divergence is constant w.r.t. $\theta$.
$KL(p|q_{\theta})=\mathbb{E}_{p(x_{1})}\left[\log p(x_{1})-\log q_{\theta}(x_{1})\right]\,$, (29)
which means that it does not enter in the gradient if we directly estimate the gradients via an MC estimator.
$\frac{dKL(p|q_{\theta})}{d\theta}$ $=-\mathbb{E}_{p}(x_{1})\left[\frac{d}{d\theta}\log q_{\theta}(x_{1})\right]$
$\approx\mathcal{G}_{ML}$ $=-\frac{1}{N}\sum_{i=1}^{N}\frac{d}{d\theta}\log q_{\theta}(x_{1}^{(i)})\,,x_{1}^{(i)}\sim p\,.$ (30)
For the exact calculation of the estimator, we decompose it fully with
$q_{\theta}(x_{1})=q_{0}(T_{\theta}^{-1}(x_{1}))\left|\det\frac{\partial T_{\theta}^{-1}(x_{1})}{\partial x_{1}}\right|\,.$ (31)
The actual calculation for the ML estimator then is
$\mathcal{G}_{ML}=-\frac{1}{N}\sum_{i=1}^{N}\frac{d}{d\theta}\left(\log q_{0}(T_{\theta}^{-1}(x_{1}))+\log\left|\frac{\partial T_{\theta}^{-1}(x_{1})}{\partial x_{1}}\right|\right)\,.$ (32)
For PG, we use Eq. 17 to directly obtain the MC estimator $\mathcal{G}_{PG}$
$\frac{d}{d\theta}KL(p_{0,\theta}|q_{0})$ $=E_{x_{1}\sim p_{1}}\left[\frac{\partial}{\partial x_{0}}\left(\log\frac{p_{0,\theta}}{q_{0}}(x_{0})\right)\frac{\partial T_{\theta}^{-1}(x_{1})}{\partial\theta}\right]$
$\approx\mathcal{G}_{PG}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial x_{0}^{(i)}}\left(\log p_{0,\theta}(x_{0}^{(i)})-\log q_{0}(x_{0}^{(i)})\right)\frac{\partial T_{\theta}^{-1}(x_{1}^{(i)})}{\partial\theta}\,,x_{1}^{(i)}\sim p\,.$ (33)
Here $x_{0}^{(i)}$ is a shorthand for $T_{\theta}^{-1}(x_{1}^{(i)})$. If we again decompose the term we expose the full calculation
$\mathcal{G}_{PG}=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial x_{0}^{(i)}}\left(\log p(T_{\theta}(x_{0}^{(i)}))+\log\left|\frac{\partial T_{\theta}(x_{0})}{\partial x_{0}}\right|-\log q_{0}(x_{0}^{(i)})\right)\frac{\partial T_{\theta}^{-1}(x_{1}^{(i)})}{\partial\theta}$ (34)
[…]
A.11.2 Variance of the estimators
Opposed to $\mathcal{G}_{ML}$ and $\mathcal{G}_{FM}$, we have nice guarantees about the variance of the PG estimator $\mathcal{G}_{PG}$ at the optimum and close to it.
We first recapitulate results from previous work and then, by using a simple example, show that $\mathcal{G}_{FM}$ does not necessarily exhibit zero variance at the optimum.
In general, the variance of the Path Gradient estimators $\mathcal{G}_{PG}$ is bounded by the squared Lipschitz constant of the term
$\left(\log p_{0,\theta}(x_{0})-\log q_{0}(x_{0})\right)=\left(\log p(x_{1})-\log q_{\theta}(x_{1})\right)\,$, (35)
see Mohamed et al. (2020) Section 5.3.2.
Thus, if the target density $p$ is not well approximated by $q_{\theta}$, the variance of the gradient estimator can be large and training with PG might not be beneficial. If $\log p$ and $\log q_{\theta}$ are close in the sense that the Lipschitz constant of their difference is small, we can assume path gradient estimators to be helpful.
Gradient estimators at the optimum
In the case of perfect approximation, i.e. $q_{\theta}(x)=p(x)\forall x\in\mathcal{X}$, the following statements about the ML estimator and the Path Gradient estimators are known. The gradients for path gradient estimators are deterministically 0, i.e. $\mathbb{E}[\mathcal{G}_{PG}]=0,\mathrm{Var}[\mathcal{G}_{PG}]=0$, while for ML the variance is generically nonzero $\mathbb{E}[\mathcal{G}_{ML}]=0,\mathrm{Var}[\mathcal{G}_{ML}]=\frac{1}{N}\mathcal{I}_{\theta}$. Where $\mathcal{I}_{\theta}$ is the Fisher Information Matrix of $p_{0,\theta}$ [2].
Why is that?
Already the review by Papamakarios et al. (2021) notes in Section 2.3.3 that the duality of the KL divergence means that fitting the model $q_{\theta}$ to the target $p$ using ML is equivalent to fitting $p_{0,\theta}$ to the base $q_{0}$ under the reverse KL. This means that the results from Vaitl (2024) directly hold. We only adapted the notation.
A.11.3 Variance of $\mathcal{G}_{FM}$ for a toy example
We do not have a term for the variance of $\mathcal{G}_{FM}$ for general settings, but we can show that it does not deterministically vanish like for $\mathcal{G}_{PG}$ via a simple example. This means better behavior for PG than for FM at the optimum, which we also verified empirically. Our assumptions aim to simplify the example as much as possible.
  • First, assume the standard loss for Flow Matching.
  • Further, assume the two densities to be the same $D$-dimensional Normal distribution
    $q_{0}=p=\mathcal{N}(0,I)$.
  • Finally, assume the CNF is the identity parametrized by a single parameter $\theta$, i.e.
    $v_{\theta}=\theta=0$.
In this example $q_{\theta}(x_{1})=q_{0}(x_{1})\cdot I$ approximates the target density $p$ perfectly and the optimal gradient estimator is 0. Yet, in this setting the variance of the estimator $\mathcal{G}_{FM}$ is non-zero:
$\mathrm{Var}[\mathcal{G}_{FM}]=\frac{8}{ND}\,$, (36)
preventing the model from staying at the optimum during training.
Proof:
The gradient estimator for the FM loss Eq (10) is
[…]
where ${\tilde{x}}_{0}^{(i)},x_{1}^{(i)}\sim\mathcal{N}(0,I)$}  [^3] Note that here ${\tilde{x}}_{0}^{(i)}$ and $x_{1}^{(i)}$ are independently sampled. Before $x_{0}^{(i)}$ was the transformed sample.
First, we break down the terms
$\mathcal{G}_{FM}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial}{\partial\theta}\frac{1}{D}\sum_{d=1}^{D}\left(v_{\theta,d}+\tilde{x}_{0,d}^{(i)}-x_{1,d}^{(i)}\right)^{2}$
$=\frac{1}{ND}\sum_{i=1}^{N}\sum_{d=1}^{D}2\left(v_{\theta,d}+\tilde{x}_{0,d}^{(i)}-x_{1,d}^{(i)}\right)\frac{\partial v_{\theta,d}}{\partial\theta}\,.$ (38)
Because $v_{\theta}$ is parametrized by $\theta$, we set in $\frac{\partial v_{\theta,d}}{\partial\theta}=1$ and $v_{\theta,d}=0$ and separate the terms
$\mathcal{G}_{FM}=\frac{1}{ND}\sum_{i}\sum_{d}2(\tilde{x}_{0,d}^{(i)}-x_{1,d}^{(i)})=\frac{2}{ND}\left(\sum_{d}\sum_{i}\tilde{x}_{0,d}^{(i)}-\sum_{d}\sum_{i}x_{1,d}^{(i)}\right).$ (39)
We can compute the distribution of $\mathcal{G}_{FM}$ by using the property
$\sum_{i=1}^{N}a_{i}\sim\mathcal{N}(0,N\sigma_{a}^{2})\text{ for }a\sim\mathcal{N}(0,\sigma_{a}^{2})\,.$ (40)
So the sum over dimensions and samples follows the normal distribution with variance $DN$:
$\sum_{d}\sum_{i}\tilde{x}_{0,d}^{(i)}\sim N(0,DN)$ (41)
and the difference has twice its variance
$\sum_{d}\sum_{i}\tilde{x}_{0,d}^{(i)}-\sum_{d}\sum_{i}x_{1,d}^{(i)}\sim\mathcal{N}(0,2DN)\,.$ (42)
The expectation of the estimator $\mathcal{G}_{FM}$ then is 0 and the variance is simply
$\mathrm{Var}[\mathcal{G}_{FM}]=\frac{4}{N^{2}D^{2}}\mathrm{Var}[(\sum_{d}\sum_{i}\tilde{x}_{0,d}^{(i)}-\sum_{d}\sum_{i}x_{1,d}^{(i)})]=\frac{4}{N^{2}D^{2}}2DN=\frac{8}{ND}\,.$ (43)
Interestingly, the derivative $\mathcal{G}_{FM}$ is invariant to re-ordering because the sums of $x_{1}$ and $\tilde{x}_{0}$ are invariant to permutation. The result thus also holds for OT-FM. This property is due to the simple assumption, the vector field $v_{\theta}$ is independent of $x_{t}$.
A.12 Pseudocode for forward KL path gradients via augmented adjoint
@@
2:  $\dot{x},\dot{\text{div}}\leftarrow\text{black\_box\_dynamics}(t,x_{t})$
@@
3:  $\nabla\log p_{1}\leftarrow\text{gradient}_{x_{1}}(\log p_{1})$ $\triangleright$ Integrate using Augmented Adjoint state method
4:  $x_{0},\nabla\log p_{0,\theta},\log|\det J|\leftarrow\text{flow.integrateAugAdjoint}(x_{1},\nabla\log p_{1}),\texttt{inverse=True})$
@@
6:  $\nabla\log q_{0}\leftarrow\text{gradient}_{x_{o}}(\log q_{0})$ $\triangleright$ Compute gradient of loss w.r.t. sample $x_{0}$
7:  $\nabla_{x_{0}}\mathcal{L}\leftarrow\frac{1}{N}\left(\nabla\log p_{0,\theta}-\nabla\log q_{0}\right)$ $\triangleright$ Backpropagate using standard Adjoint state method
8:  $\text{path gradients}\leftarrow\text{gradient}_{\theta}\left(x_{0}\cdot\texttt{detach}(\nabla_{x_{0}}\mathcal{L})]\right)$
9:end function
A.13 Code libraries